Chair: Eloise

Next meeting chair: Prad

Attended: Jasper, Simin, Jamie,

Apologies:

General Issues

  • Hobart 26m out of action till ~end of year
  • Refer to new wiki for all setups, make notes of anything missing/that's needed to be added
  • Checklists will be in use again, please be sure to do these
  • slogit not able to get sched file off of BKG server, notes on wiki if this happens again. Let Jamie know if you encounter any issues when downloading sched file

AUM036

Hobart 12m

  • 2305 - S-band, X-pol frequency settings corrected (JMc).
  • 0624 - Sampler stats report as bad. ADB/Core3H reset started at 0640 (JMc)
  • Unsuccessful, attempting reboot but apparent issues with UPS. (JMc) - Fiddled with setting on NUT UPS server (notes are in wiki)